OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5224 — tou59-1-3k
Scheduled Maintenance Report for Web Cloud
Completed
We detected a problem of optical on tou59-1-3k :

%SFF8472-5-THRESHOLD_VIOLATION: Te0/1: Rx power low alarm; Operating value: -16.0 dBm, Threshold value: -15.2 dBm.

We will intervene tomorrow morning at 10h00. An interruption of about 1 hour is predicted. The customers on the NRA of Roncq will be impacted during this maintenance.



Update(s):

Date: 2011-03-11 12:24:33 UTC
We have just cleaned out the optic bloc in addition to the fibre, the optic power supply of reception is back to normal. The NRA of roncq is up again.

Optical Optical
Temperature Voltage Tx Power Rx Power
Port (Celsius) (Volts) (dBm) (dBm)
--------- ----------- ------- -------- --------
Te0/1 44.8 3.38 -1.6 -11.2

Date: 2011-03-11 12:21:52 UTC
We will ask for the intervention. The NRA of roncq will be available during the tests and the optic replacement.
Posted Mar 10, 2011 - 15:48 UTC